# On the controller form for linear hyperbolic MIMO systems with dynamic boundary conditions

### v1

@@ IV-D3 Generalised polynomial long division
Let $x(\sigma,\partial_{t}),y(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{r}},\;y(\sigma,\partial_{t})\neq 0$ be generalised polynomials with rational coefficients where $\deg_{\sigma}x(\sigma)\geq\deg_{\sigma}y(\sigma)$. Then there exist $q(\sigma,\partial_{t}),r(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{r}}$ such that
$x(\sigma,\partial_{t})=q(\sigma,\partial_{t})y(\sigma,\partial_{t})+r(\sigma,s)$ (27)
with $\deg_{\sigma}r(\sigma,\partial_{t})<\deg_{\sigma}y(\sigma,\partial_{t})$ as well as $\deg^{+}_{\sigma}r(\sigma,\partial_{t})<\deg^{+}_{\sigma}y(\sigma,\partial_{t})$ and $\deg^{-}_{\sigma}r(\sigma,\partial_{t})>\deg^{-}_{\sigma}y(\sigma,\partial_{t})$.
@@
As the following steps are valid for arbitrary coefficients, the explicit mention of $\partial_{t}$ will be omitted. In the general case we have $\deg_{\sigma}x(\sigma)>\deg_{\sigma}y(\sigma)$ as well as $\deg^{+}_{\sigma}x(\sigma)>\deg^{+}_{\sigma}y(\sigma)$, $\deg^{-}_{\sigma}x(\sigma)<\deg^{-}_{\sigma}y(\sigma)$ or both. Since the exponents of the polynomials $x(\sigma,\partial_{t})=\sum_{i=1}^{n}a_{i}(\partial_{t})\sigma^{\alpha_{i}}$, $y(\sigma,\partial_{t})=\sum_{i=1}^{m}b_{i}(\partial_{t})\sigma^{\beta_{i}}$ are ordered, we use
$q(\sigma)=\begin{cases}\frac{a_{1}\sigma^{\alpha_{1}}}{b_{1}\sigma^{\beta_{1}}},&\text{ if }\deg_{\sigma}^{+}x(\sigma)>\deg^{+}_{\sigma}y(\sigma)\\ \frac{a_{n}\sigma^{\alpha_{n}}}{b_{m}\sigma^{\beta_{m}}},&\text{ if }\deg^{-}_{\sigma}x(\sigma)<\deg^{-}_{\sigma}y(\sigma),\end{cases}$ (28)
@@
$\begin{split}r(\sigma)=x(\sigma)-q(\sigma)y(\sigma)=\sum\limits_{i=2}^{n}a_{i}\sigma^{\alpha_{i}}-\frac{a_{1}\sigma^{\alpha_{1}}}{b_{1}\sigma^{\beta_{1}}}\sum\limits_{i=2}^{m}b_{i}\sigma^{\beta_{i}},\end{split}$
@@
as well as $\deg^{-}_{\sigma}r(\sigma)=\min\{\alpha_{n},\alpha_{1}-(\beta_{1}-\beta_{m})\}$ in which we identify $\deg_{\sigma}y(\sigma)=\beta_{1}-\beta_{m}$ and use $\alpha_{1}=\deg_{\sigma}x(\sigma)+\alpha_{n}$ as well as $\kappa=\deg_{\sigma}x(\sigma)-\deg_{\sigma}y(\sigma)\geq 0$ to obtain
$\begin{split}\deg^{-}_{\sigma}r(\sigma)=\min\{&\alpha_{n},\alpha_{n}+\kappa\}=\alpha_{n}=\deg^{-}_{\sigma}x(\sigma).\end{split}$
This shows that the predictions in $x(\sigma)$ can be reduced without affecting the delays. Regarding the overall degree, we have $\deg_{\sigma}r(\sigma)=\max\{\alpha_{2}-\alpha_{n},\alpha_{1}-\alpha_{n}-(\beta_{1}-\beta_{2})\}$ which yields $\deg_{\sigma}r(\sigma)<\deg_{\sigma}x(\sigma)$. Hence, $x(\sigma)$ can be reduced in a manner such that the iteration
@@
with $r_{0}(\sigma)=x(\sigma)$ and $q_{k}(\sigma)=\frac{c_{k,1}\sigma^{\gamma_{k,1}}}{b_{1}\sigma^{\beta_{1}}}$ by $r_{k}(\sigma)=\sum_{i=1}^{n}c_{k,i}\sigma^{\gamma_{k,i}}$ will for some $k^{+}$ terminate with $\deg^{+}_{\sigma}r_{k^{+}}(\sigma)<\deg^{+}_{\sigma}y(\sigma)$ since $x(\sigma)$ only has a finite amount of summands.
Now, (if required) the second branch of (28) is taken and similar computations yield that restarting the iteration with $r_{0}(\sigma)=r_{k^{+}}(\sigma)$ and choosing $q_{k}(\sigma)=\frac{c_{k,1}\sigma^{\gamma_{k,1}}}{b_{m}\sigma^{\beta_{m}}}$ will yield $\deg^{-}_{\sigma}r_{k^{-}}(\sigma)>\deg^{-}_{\sigma}y(\sigma)$ after $k^{-}$ more steps. Summing up all $q_{k}(\sigma)$ from both iterations then yields the resulting “quotient” $q(\sigma)$ with common denominator $d=b_{1}^{k^{+}}b_{m}^{k^{-}}$. ∎
With Lemma 2, the shift degrees are taken care of. However, the coefficients of $q(\sigma,\partial_{t}),r(\sigma,\partial_{t})$ are rational expressions in $\partial_{t}$. Furthermore, in the analytic setting, their denominators can be interpreted as functions that their numerator polynomials are applied to. This results in piecewise continuous functions which do not necessarily have compact support. In other words: They generally do not correspond to differential or (distributed) delay operators with finite shift amplitude from $\mathcal{E}^{\prime}$. However, thanks to the Palay-Wiener theorem, if their Laplace transforms are entire functions, they indeed are such operators, cf. [11, Thm. 7.3.1].
@@
Let $x(\sigma,\partial_{t}),y(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{p}},\;y(\sigma,\partial_{t})\neq 0$ be generalised polynomials with polynomial coefficients where $\deg_{\sigma}x(\sigma)\geq\deg_{\sigma}y(\sigma)$, as well as $\deg_{\partial_{t}}x(\sigma)\leq\deg_{\partial_{t}}y(\sigma)$. Then there exist $q^{\star}(\sigma,\partial_{t}),r^{\star}(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{w}}=\mathcal{R}_{\mathrm{r}}\cap\mathcal{E}^{\prime}$ such that
$x(\sigma,\partial_{t})=q^{\star}(\sigma,\partial_{t})y(\sigma,\partial_{t})+r^{\star}(\sigma,s)$ (29)
with $\deg_{\sigma}r^{\star}(\sigma,\partial_{t})=\deg_{\sigma}y(\sigma,\partial_{t})$, $\deg^{+}_{\sigma}r^{\star}(\sigma,\partial_{t})=\deg^{+}_{\sigma}y(\sigma,\partial_{t})$ and $\deg^{-}_{\sigma}r^{\star}(\sigma,\partial_{t})=\deg^{-}_{\sigma}y(\sigma,\partial_{t})$.
@@
This proof is heavily inspired by [12, 9] and only adapted to the current setup. First, we obtain $q(\sigma,\partial_{t}),r(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{r}}$ by means of Lemma 2 and extract its normalised common denominator $d(\partial_{t})\in\mathbb{R}[\partial_{t}]$ such that $q(\sigma,\partial_{t})=\frac{\tilde{q}(\sigma,\partial_{t})}{d(\partial_{t})}$ with $\tilde{q}(\sigma,\partial_{t})\in\mathcal{R}_{\mathrm{p}}$. Next, the factorisation $d(\partial_{t})=\prod_{k=1}^{D}(\partial_{t}-s_{k})$ with the roots $s_{k}$ and $D=\deg d(\partial_{t})$ are used together with $\hat{q}_{0}(s)=\mathcal{L}\{\tilde{q}(\sigma,\partial_{t})\}$ in the iteration
@@
which introduces a nominator zero for every pole from $q(\sigma,\partial_{t})$ it adds. Thus, $q_{D}(s)$ obtained after $D$ iterations is an entire function on $\mathbb{C}$ and therefore $q^{\star}(\sigma,\partial_{t})=\mathcal{L}^{-1}\{q_{D}(s)\}\in\mathcal{R}_{\mathrm{w}}$ by [11, Thm. 7.3.1].
Now, due to the structure of (30) we can justify the separation $q^{\star}(\sigma,\partial_{t})=q(\sigma,\partial_{t})+p(\partial_{t})$ with $p(\partial_{t})\in\mathbb{R}(\partial_{t})$. Therefore, we obtain the degree conditions from Lemma 2 with equal signs for the new remainder $r^{\star}(\sigma,\partial_{t})$. ∎
IV-D4 Algorithm
Equipped with a suitable way to reduce the degrees of its elements by means of the GPLD from Lemma 3, we now reduce the degrees in the entire matrix $\mathbf{H}(\sigma,\partial_{t})$ by applying a series of row transformations.
Algorithm 1 Shift reduction in $\mathbf{H}(\sigma,\partial_{t})$
@@ IV-D4 Algorithm
1: $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]function compute_transform($\mathbf{X},j$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
@@
5:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $q^{\star}_{i,j}\leftarrow$ GPLD($x=\mathbf{X}[i,j],y=\mathbf{X}[j,j]$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
@@
9: $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]procedure reduce_shifts($\mathbf{H}$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
@@
10:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\bar{}\mathbf{H}\leftarrow\mathbf{H}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
11:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]for $i\in\{2,\dotsc,{n^{-}}\}$ do $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
12:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]for $j\in\{1,\dotsc,{n^{-}}-1\}$ do $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
13:     $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\mathbf{L}_{c}\leftarrow$ Compute transform($\bar{}\mathbf{H},j$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
@@
14:     $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\bar{}\mathbf{H}\leftarrow\mathbf{L}_{c}\mathbf{H}$, $\mathbf{L}\leftarrow\mathbf{L}_{c}\mathbf{L}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
@@
15:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] return $\bar{}\mathbf{H}$$\mathbf{L}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]

### v3

@@ IV-D3 Quasipolynomial long division on $\mathcal{R}_{\mathrm{r}}$
Let $x(\sigma,{\mathrm{d}_{t}}),y(\sigma,{\mathrm{d}_{t}})\in\mathcal{R}_{\mathrm{r}},\;y(\sigma,{\mathrm{d}_{t}})\neq 0$ be quasipolynomials in $\sigma$ with rational coefficients in $\mathrm{d}_{t}$} and $\deg_{\sigma}x(\sigma)\geq\deg_{\sigma}y(\sigma)$ as well as $\deg_{{\mathrm{d}_{t}}}x(\sigma)\leq\deg_{{\mathrm{d}_{t}}}y(\sigma)$. Then there exist $q(\sigma,{\mathrm{d}_{t}}),r(\sigma,{\mathrm{d}_{t}})\in\mathcal{R}_{\mathrm{r}}$ such that
$x(\sigma,{\mathrm{d}_{t}})=q(\sigma,{\mathrm{d}_{t}})y(\sigma,{\mathrm{d}_{t}})+r(\sigma,s)$ (27)
with $\deg_{\sigma}r(\sigma,{\mathrm{d}_{t}})<\deg_{\sigma}y(\sigma,{\mathrm{d}_{t}})$, $\deg^{+}_{\sigma}r(\sigma,{\mathrm{d}_{t}})<\deg^{+}_{\sigma}y(\sigma,{\mathrm{d}_{t}})$ and $\deg^{-}_{\sigma}r(\sigma,{\mathrm{d}_{t}})>\deg^{-}_{\sigma}y(\sigma,{\mathrm{d}_{t}})$.
@@
As the following steps are valid for arbitrary coefficients, the explicit mention of $\mathrm{d}_{t}$} will be omitted. In the general case we have $\deg_{\sigma}x(\sigma)>\deg_{\sigma}y(\sigma)$ as well as $\deg^{+}_{\sigma}x(\sigma)>\deg^{+}_{\sigma}y(\sigma)$, $\deg^{-}_{\sigma}x(\sigma)<\deg^{-}_{\sigma}y(\sigma)$ or both. Since the exponents of $x(\sigma,{\mathrm{d}_{t}})=\sum_{i=1}^{u}a_{i}({\mathrm{d}_{t}})\sigma^{\alpha_{i}}$, $y(\sigma,{\mathrm{d}_{t}})=\sum_{i=1}^{v}b_{i}({\mathrm{d}_{t}})\sigma^{\beta_{i}}$ are ordered, we use
$q(\sigma)=\begin{cases}\frac{a_{1}}{b_{1}}\sigma^{\alpha_{1}-\beta_{1}},&\text{ if }\deg_{\sigma}^{+}x(\sigma)>\deg^{+}_{\sigma}y(\sigma)\\ \frac{a_{u}}{b_{v}}\sigma^{\alpha_{u}-\beta_{v}},&\text{ if }\deg^{-}_{\sigma}x(\sigma)<\deg^{-}_{\sigma}y(\sigma),\end{cases}$ (28)
@@
$\begin{split}r(\sigma)=x(\sigma)-q(\sigma)y(\sigma)=\sum\limits_{i=2}^{u}a_{i}\sigma^{\alpha_{i}}\!-\!\frac{a_{1}}{b_{1}}\sigma^{\alpha_{1}-\beta_{1}}\!\sum\limits_{i=2}^{v}b_{i}\sigma^{\beta_{i}}\end{split}$
@@
as well as $\deg^{-}_{\sigma}r(\sigma)=\min\{\alpha_{u},\alpha_{1}-(\beta_{1}-\beta_{v})\}$ in which we identify $\deg_{\sigma}y(\sigma)=\beta_{1}-\beta_{v}$ and use $\alpha_{1}=\deg_{\sigma}x(\sigma)+\alpha_{u}$ as well as $\kappa=\deg_{\sigma}x(\sigma)-\deg_{\sigma}y(\sigma)\geq 0$ to obtain
$\begin{split}\deg^{-}_{\sigma}r(\sigma)=\min\{&\alpha_{u},\alpha_{u}+\kappa\}=\alpha_{u}=\deg^{-}_{\sigma}x(\sigma).\end{split}$
This shows that the predictions in $x(\sigma)$ can be reduced without affecting the delays. Regarding the overall degree, we have $\deg_{\sigma}r(\sigma)=\max\{\alpha_{2}-\alpha_{u},\alpha_{1}-\alpha_{u}-(\beta_{1}-\beta_{2})\}$ which yields $\deg_{\sigma}r(\sigma)<\deg_{\sigma}x(\sigma)$. Hence, $x(\sigma)$ can be reduced in a manner such that the iteration
@@
with $r_{0}(\sigma)=x(\sigma)$ and $q_{k}(\sigma)=\frac{c_{k,1}}{b_{1}}\sigma^{\gamma_{k,1}-\beta_{1}}$ by $r_{k}(\sigma)=\sum_{i=1}^{w_{k}}c_{k,i}\sigma^{\gamma_{k,i}}$ will for some $k^{+}$ terminate with $\deg^{+}_{\sigma}r_{k^{+}}(\sigma)<\deg^{+}_{\sigma}y(\sigma)$ since $x(\sigma)$ only has a finite number of summands.
If required, the second branch of (28) is then taken and similar computations yield that restarting the iteration with $r_{0}(\sigma)=r_{k^{+}}(\sigma)$ and choosing $q_{k}(\sigma)=\frac{c_{k,1}}{b_{v}}\sigma^{\gamma_{k,1}-\beta_{v}}$ will yield $\deg^{-}_{\sigma}r_{k^{-}}(\sigma)>\deg^{-}_{\sigma}y(\sigma)$ after $k^{-}$ more steps. Summing up all $q_{k}(\sigma)$ from both iterations then yields the resulting “quotient” $q(\sigma)$ with denominator $d=b_{1}^{k^{+}}b_{v}^{k^{-}}$. ∎
IV-D4 Quasipolynomial long division on $\mathcal{R}_{\mathrm{e}}$
With Lemma 2, the shift degrees have been taken care of. However, the coefficients of $q(\sigma,{\mathrm{d}_{t}})$ and $r(\sigma,{\mathrm{d}_{t}})$ still are rational expressions in $\mathrm{d}_{t}$}. Since, in the analytic setting, their denominators can be interpreted as functions that their numerator polynomials are applied to, the results are piecewise continuous functions which do not necessarily have compact support. In other words: They generally do not correspond to differential or (distributed) delay operators with finite shift amplitude from $\mathcal{E}^{\prime}$ and can therefore not be used to reduce the shifts as they would introduce indefinite integrals.
Nevertheless, thanks to the Paley-Wiener theorem [21, Thm. 7.3.1], if their Laplace transforms are entire functions, they indeed are such operators. For example, take $S_{z}=\frac{1}{2}\frac{\sigma^{z}-\sigma^{-z}}{{\mathrm{d}_{t}}}\in\mathcal{R}_{\mathrm{r}}$ from the introduction, whose Laplace transform $\mathcal{L}\{S_{z}\}(s)=\frac{1}{2s}\left(e^{sz}-e^{-sz}\right)$ appears to possess a pole at $s=0$. However, since its numerator introduces a zero at the same location, this pole can be lifted rendering $\mathcal{L}\{S_{z}\}(s)$ an entire function and thus yielding $S_{z}\in\mathcal{E}^{\prime}$ (which was already clear from (8)). Thus, the idea is to introduce numerator zeros to $q(\sigma,{\mathrm{d}_{t}})$ in form of a correction term $p({\mathrm{d}_{t}})$ such that a modified $q^{\star}(\sigma,{\mathrm{d}_{t}})=q(\sigma,{\mathrm{d}_{t}})+p({\mathrm{d}_{t}})$ belongs to $\mathcal{R}_{\mathrm{e}}=\mathcal{R}_{\mathrm{r}}\cap\mathcal{E}^{\prime}$ and is therefore admissible.
@@ IV-D4 Quasipolynomial long division on $\mathcal{R}_{\mathrm{e}}$
Let $x(\sigma,{\mathrm{d}_{t}}),y(\sigma,{\mathrm{d}_{t}})\in\mathcal{R}_{\mathrm{e}},\;y(\sigma,{\mathrm{d}_{t}})\neq 0$ be operators from $\mathcal{E}^{\prime}$, written as quasipolynomials in $\sigma$ with rational coefficients in $\mathrm{d}_{t}$} and $\deg_{\sigma}x(\sigma)\geq\deg_{\sigma}y(\sigma)$, as well as $\deg_{{\mathrm{d}_{t}}}x(\sigma)\leq\deg_{{\mathrm{d}_{t}}}y(\sigma)$. Then there exist $q^{\star}(\sigma,{\mathrm{d}_{t}}),r^{\star}(\sigma,{\mathrm{d}_{t}})\in\mathcal{R}_{\mathrm{e}}$ such that
$x(\sigma,{\mathrm{d}_{t}})=q^{\star}(\sigma,{\mathrm{d}_{t}})y(\sigma,{\mathrm{d}_{t}})+r^{\star}(\sigma,s)$ (29)
with $\deg_{\sigma}r^{\star}(\sigma,{\mathrm{d}_{t}})=\deg_{\sigma}y(\sigma,{\mathrm{d}_{t}})$, $\deg^{+}_{\sigma}r^{\star}(\sigma,{\mathrm{d}_{t}})=\deg^{+}_{\sigma}y(\sigma,{\mathrm{d}_{t}})$ and $\deg^{-}_{\sigma}r^{\star}(\sigma,{\mathrm{d}_{t}})=\deg^{-}_{\sigma}y(\sigma,{\mathrm{d}_{t}})$.
@@
This proof is heavily inspired by [9, 10] and only adapted to the current setup. First, we obtain $q(\sigma,{\mathrm{d}_{t}}),r(\sigma,{\mathrm{d}_{t}})\in\mathcal{R}_{\mathrm{r}}$ by means of Lemma 2 and extract its normalised common denominator $d({\mathrm{d}_{t}})\in\mathbb{R}[{\mathrm{d}_{t}}]$ such that $q(\sigma,{\mathrm{d}_{t}})=\frac{\tilde{q}(\sigma,{\mathrm{d}_{t}})}{d({\mathrm{d}_{t}})}$ with $\tilde{q}(\sigma,{\mathrm{d}_{t}})\in\mathcal{R}_{\mathrm{p}}$. Next, the factorisation $d({\mathrm{d}_{t}})=\prod_{k=1}^{D}({\mathrm{d}_{t}}-s_{k})$ with the roots $s_{k}$ and $D=\deg d({\mathrm{d}_{t}})$ are used together with $\hat{q}_{0}(s)=\mathcal{L}\{\tilde{q}(\sigma,{\mathrm{d}_{t}})\}$ in the iteration
@@
which introduces a numerator zero for every pole from $q(\sigma,{\mathrm{d}_{t}})$ it adds. Thus, $q_{D}(s)$ obtained after $D$ iterations is an entire function on $\mathbb{C}$ and therefore $q^{\star}(\sigma,{\mathrm{d}_{t}})=\mathcal{L}^{-1}\{q_{D}(s)\}\in\mathcal{R}_{\mathrm{e}}$ by [21, Thm. 7.3.1].
Now, due to the structure of (30) we can justify the separation $q^{\star}(\sigma,{\mathrm{d}_{t}})=q(\sigma,{\mathrm{d}_{t}})+p({\mathrm{d}_{t}})$ with $p({\mathrm{d}_{t}})\in\mathbb{R}({\mathrm{d}_{t}})$. Therefore, we obtain the degree conditions from Lemma 2 with equal signs for the new remainder $r^{\star}(\sigma,{\mathrm{d}_{t}})$. ∎
IV-D5 Algorithm
Equipped with a suitable way to reduce the degrees of its elements by means of the QPLD from Lemma 3, we now reduce the degrees in the entire matrix $\mathbf{G}(\sigma,{\mathrm{d}_{t}})$ by applying a series of row transformations.
Algorithm 1 Shift reduction in $\mathbf{G}(\sigma,{\mathrm{d}_{t}})$
@@ IV-D5 Algorithm
1: $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]function build_transform($\mathbf{X},j$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
@@
5:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $q^{\star}_{i,j}\leftarrow$ qpld($x=\mathbf{X}[i,j],y=\mathbf{X}[j,j]$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
@@
9: $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]procedure reduce_shifts($\mathbf{G}$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
@@
10:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\bar{}\mathbf{G}\leftarrow\mathbf{G},\;\mathbf{L}\leftarrow\mathbf{I}_{{n^{-}}}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
@@
11:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]while $\sum_{i=0}^{{n^{-}}}\operatorname{coldeg}_{i}\mathbf{\bar{G}}>\tau_{\Sigma}$ do $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
@@
12:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]if $\deg_{\sigma}\bar{g}_{ii}\neq\deg_{\sigma}\bar{g}_{jj}\;\forall i,j\in\{1,\dotsc,{n^{-}}\}$ then $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
13:     $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]for $j\in\{1,\dotsc,{n^{-}}-1\}$ do $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
14:      $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\mathbf{L}_{c}\leftarrow$ build_transform($\bar{}\mathbf{G},j$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
15:      $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\bar{}\mathbf{G}\leftarrow\mathbf{L}_{c}\mathbf{G}$, $\mathbf{L}\leftarrow\mathbf{L}_{c}\mathbf{L}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
16:   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]else $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
17:     $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\mathbf{L}_{c}\leftarrow$ mqpld($\mathbf{\bar{G}}$) $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
18:     $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] $\bar{}\mathbf{G}\leftarrow\mathbf{L}_{c}\mathbf{G}$, $\mathbf{L}\leftarrow\mathbf{L}_{c}\mathbf{L}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]   $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]
19:  $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]] return $\bar{}\mathbf{G}$$\mathbf{L}$ $[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]$[$ Image Discarded Due To ‘/tikz/external/mode=list and make’ ]]